# Learning responsibility allocations for multi-agent interactions: A differentiable optimization approach with control barrier functions

### v1

@@ VI Experiments on synthetic dataset
For each system, we randomly sample 128 state and desired control pairs and then, for some choice of $\gamma$, we apply the responsibility-aware CBF filter described in Prob. 3, and then add zero-mean, Gaussian noise (variance of 0.1)0.1)0.1 ) to the solutions. Then we execute Alg. 1 to learn $\gamma$. As shown in Figs. 3(a) and 3(b), for a random initial guess, our estimated responsibility allocation $\gamma$ quickly converges to the ground truth value, even if the ground truth is time-varying (see Fig. 3(a)). We use a mini-batch size of 8 and a step size of $\delta=0.005$ and $\delta=0.05$ for the 2-agent and 6-agent problems, respectively. Additionally, Fig. 3(c) presents computation times for calculating the loss and gradient for Prob. 4. We see that our approach has great potential for real-time applications, such as estimating responsibility allocation online, which would only require small batches of data that are streaming in real-time. The experiments were performed on a MacBook Pro with a M2 processor.
@@ VII Experiments on traffic-weaving interaction
Figure 7: Comparison of $\gamma$ landscapes when learned using different subsets of the traffic weaving dataset. In the displayed scenario, the red car is moving faster than the blue car, and the black arrows represent their desired controls as they move to their goals.
@@ VII-A1 Traffic-weaving dynamics and CBF construction
We model each agent with double integrator dynamics, and consider their relative state, $\mathbf{r}=[r_{\mathrm{lon}},r_{\mathrm{lat}},\dot{r}_{\mathrm{lon}},\dot{r}_{\mathrm{lat}}]=\mathbf{x}_{2}-\mathbf{x}_{1}$ where $\mathbf{x}_{i}=[x_{\mathrm{lon}},x_{\mathrm{lat}},\dot{x}_{\mathrm{lon}},\dot{x}_{\mathrm{lat}}]^{T}$ is the state of agent $i$, and $\mathrm{lon}$ and $\mathrm{lat}$ denotes the longitudinal and lateral directions respectively. With this relative double integrator dynamics, we select a CBF $b(\mathbf{r})=\frac{r_{\mathrm{lon}}^{2}}{a_{1}^{2}}+\frac{r_{\mathrm{lat}}^{2}}{a_{2}^{2}}-1$. We set $a_{1}=9.22$ and $a_{2}=1.76$ based on the car’s sizes. Due to the double integrator being relative degree two, we employ a high-order CBF (see Remark 3).
@@
As mentioned previously, the traffic-weaving dataset consists of diverse behaviors owing to the various driving styles of the driving participants. As such, we consider four different subsets of the dataset to perform our analysis. (D1) Single traffic-weaving trajectory. (D2) Trajectories where both cars start side-by-side at equal initial longitudinal speed. (D3) Trajectories where one car starting behind but traveling faster overtakes the other car. (D4) Full dataset.
@@
We parameterize the responsibility allocation function as a neural network $\phi$ that takes in the relative state $\mathbf{r}$ and outputs the responsibility allocation. Our $\phi(\mathbf{r})$ function is a small multi-layer perception with three linear hidden layers, hidden layer size of sixteen, and $\tanh$ activation functions, whose final output is passed through the function in (8) to enforce symmetry. Our setup used JAX [25] and Equinox [29]. We trained on each dataset for $3000$ epochs with batch sizes of $16$ for the single weaving trajectory and $256$ for the other datasets, and used the ADAM optimizer [30] with a learning rate of $10^{-3}$. Our loss function $\Delta$ from Prob. 4 is the Huber loss, and the regularization/slack variables $\beta_{1}$ and $\beta_{2}$ from Prob. 1 are tuned to $0.1$ and $600$ respectively.
@@ VII-B Training on a single trajectory (D1)
We first apply our learning method to the D1 dataset for a sanity check and validate our intuition. In this single traffic-weaving trajectory (similar to Fig. 4), the car in the upper lane starts $2$m/s faster and $3$m behind the lower car, before overtaking the lower car and switching lanes first.
@@ VII-C Training on data with same initial conditions (D2)
We apply our learning method to the D2 dataset to see what happens when there are diverse and contradictory behaviors within the dataset. As the initial conditions place the cars on equal footing at the start, there is ambiguity as to who passes whom, resulting in multimodal behavior within the dataset. Learning responsibility allocations in this setting is substantially more difficult; as shown by Fig. 7b. The learned responsibility allocation is constant—regardless of their relative state, both agents are predicted to be equally responsible. This performance drop under multimodal interactions highlights a limitation in our approach, necessitating future work on a probabilistic extension.
@@ VII-D Training on data when rear, faster car overtakes (D3)
We apply our learning method to the D3 dataset to investigate if it can successfully capture the trend that the slower car yields to the faster car. In Fig. 7c, we see that with the same longitudinal position, the faster car has a much larger $\gamma$ value, meaning they are less responsible for maintaining safety, consistent with behavior in the dataset where the slower car yields to the faster car.
@@ VII-E Training on the whole dataset (D4)
Finally, we apply our learning methods to the entire traffic-weaving dataset (D4). Fig. 7d shows that it learned a responsibility allocation model similar to the model trained on D3. Despite D4 containing more diverse trajectories, there is a bias towards cars starting behind but traveling faster to overtake rather than yield to the slower but front car [28]. We see that our model can capture the bias in the dataset.
@@ VIII Conclusions and future work
In this paper, we introduced a novel responsibility allocation framework that distills ill-defined social norms underpinning human interactions into an interpretable quantity. We presented a mathematical formulation of responsibility rooted in Control Barrier Functions and a computationally efficient approach to learn responsibility allocations from data via differentiable optimization. We also proposed an approach to learning symmetric responsibility allocations which can help with data efficiency. We demonstrated the efficacy of our algorithm on synthetic and human interaction data, and results indicate that we can gain interpretable insights into the social dynamics of multi-agent interactions. For future work, we seek to (i) investigate principled ways to construct the desired control policy, e.g. via learned generative models, (ii) develop a probabilistic extension to this framework to account for multimodal interactions, and (iii) examine the use of responsibility allocations for guiding robot policy construction.

### v2

@@ VI Experiments on synthetic dataset
For each system, we randomly sample 128 state and desired control pairs and then, for some choice of $\gamma$, we apply the responsibility-aware CBF filter described in Prob. 3, and then add zero-mean, Gaussian noise (variance of $0.1)$ to the solutions. Then we execute Alg. 1 to learn $\gamma$. As shown in Figs. 3(a) and 3(b), for a random initial guess, our estimated responsibility allocation $\gamma$ quickly converges to the ground truth value, even if the ground truth is time-varying (see Fig. 3(a)). We use a mini-batch size of 8 and a step size of $\delta=0.005$ and $\delta=0.05$ for the 2-agent and 6-agent problems, respectively. Additionally, Fig. 3(c) presents computation times for calculating the loss and gradient for Prob. 4. We see that our approach has great potential for real-time applications, such as estimating responsibility allocation online, which would only require small batches of data that are streaming in real-time. The experiments were performed on a MacBook Pro with a M2 processor.
@@ VII Experiments on traffic-weaving interaction
Figure 7: Comparison of $\gamma$ landscapes when learned using different subsets of the traffic weaving dataset (D2, D3, D4; see Sec. VII-A3 for details). In this scenario, red is moving faster than blue, with the black arrows representing their desired goal-reaching controls.
@@ VII-A1 Traffic-weaving dynamics and CBF construction
We model each agent with double integrator dynamics, and consider their relative state, $\mathbf{r}=[r_{\mathrm{lon}},r_{\mathrm{lat}},\dot{r}_{\mathrm{lon}},\dot{r}_{\mathrm{lat}}]=\mathbf{x}_{2}-\mathbf{x}_{1}$ where $\mathbf{x}_{i}=[x_{\mathrm{lon}},x_{\mathrm{lat}},\dot{x}_{\mathrm{lon}},\dot{x}_{\mathrm{lat}}]^{T}$ is the state of agent $i$, and $\mathrm{lon}$ and $\mathrm{lat}$ denotes the longitudinal and lateral directions respectively. Using relative states reduces the dynamical system’s dimensionality. With these relative double integrator dynamics, we select a CBF $b(\mathbf{r})=\frac{r_{\mathrm{lon}}^{2}}{a_{1}^{2}}+\frac{r_{\mathrm{lat}}^{2}}{a_{2}^{2}}-1$. We set $a_{1}=9.22$ and $a_{2}=1.76$ based on the cars’ sizes. Due to the double integrator being relative degree two, we employ a high-order CBF (see Remark 3).
@@
As mentioned previously, the traffic-weaving dataset consists of diverse behaviors owing to the various driving styles of the driving participants. As such, we consider four different subsets of the dataset to perform our analysis. (D1) Single traffic-weaving trajectory. (D2) Trajectories where both cars start side-by-side at equal initial longitudinal speed. (D3) Trajectories where one car starting behind but traveling faster overtakes the other car. (D4) Whole dataset.
@@
We parameterize the responsibility allocation function as a neural network $\phi$ that takes in the relative state $\mathbf{r}$ and outputs the responsibility allocation. Our $\phi(\mathbf{r})$ function is a small multi-layer perception with three linear hidden layers, hidden layer size of sixteen, and $\tanh$ activation functions, whose final output is passed through the function in (8) to enforce symmetry. Our setup used JAX [7] and Equinox [11]. We trained on each dataset for $3000$ epochs with batch sizes of $16$ for the single weaving trajectory and $256$ for the other datasets, and used the ADAM optimizer [12] with a learning rate of $10^{-3}$. Our loss function $\Delta$ from Prob. 4 is the Huber loss, and the regularization/slack coefficients $\beta_{1}$ and $\beta_{2}$ from Prob. 1 are tuned to $0.1$ and $600$ respectively.
@@ VII-B Training on a single trajectory (D1)
We first apply our learning method to the D1 dataset for a sanity check and validate our intuition. In this single traffic-weaving trajectory (similar to Fig. 4), the car in the upper lane starts 22m/s faster and 33m behind the lower car, before overtaking the lower car and switching lanes first.
@@ VII-C Training on data with same initial conditions (D2)
We apply our learning method to the D2 dataset to see what happens when the dataset contains diverse and contradictory behaviors. As the initial conditions place the cars on equal footing at the start, there is ambiguity as to who passes whom, resulting in multimodal behavior within the dataset. Learning responsibility allocations in this setting is substantially more difficult; as shown by Fig. 7: Same initial states. The learned responsibility allocation is constant—regardless of their relative state, both agents are predicted to be equally responsible. This performance drop under multimodal interactions highlights a limitation in our approach, necessitating future work on a probabilistic extension.
@@ VII-D Training on data when rear, faster car overtakes (D3)
We apply our learning method to the D3 dataset to investigate if it can successfully capture the trend that the slower car yields to the faster car. In Fig. 7: Rear car overtakes, we see that with the same longitudinal position, the faster car has a much larger $\gamma$ value, meaning they are less responsible for maintaining safety, consistent with behavior in the dataset where the slower car yields to the faster car.
@@ VII-E Training on the whole dataset (D4)
Finally, we apply our learning methods to the entire traffic-weaving dataset (D4). Fig. 7: Whole dataset shows that it learned a responsibility allocation model similar to the model trained on D3. Despite D4 containing more diverse trajectories, there is a bias towards cars starting behind but traveling faster to overtake rather than yield to the slower but front car [21]. Clearly, our model captures this bias.
@@ VIII Conclusions and future work
In this paper, we introduced a novel responsibility allocation framework that distills ill-defined social norms underpinning human interactions into an interpretable quantity. We presented a mathematical formulation of responsibility rooted in Control Barrier Functions and a computationally efficient approach to learn responsibility allocations from data via differentiable optimization. We also proposed an approach to learning symmetric responsibility allocations, improving data efficiency. We showed the efficacy of our algorithm on synthetic and human interaction data, and results indicate we can gain interpretable insights into the social dynamics of multi-agent interactions. For future work, we will (i) investigate principled ways to construct desired control policies, e.g. via learned generative models, (ii) develop a probabilistic extension to this framework to account for multimodal interactions, and (iii) examine using responsibility allocations for guiding robot policy construction.